OpenStudy (amber15):

which bottle of grapefruit juice has the lower unit price 64 oz for 3.96 or 48 oz for 2.96 my answer is 64 oz some one tell me if i am correct

OpenStudy (hunus):

That's correct

OpenStudy (amber15):

thank you

OpenStudy (hunus):

Yup :)

OpenStudy (amber15):

i hope im right cause this is what i figured out : A=3.96/64=0.0618 price/oz. B=2.96/48=0.0616 price/oz. We can see that B has the lower unit price

OpenStudy (amber15):

so it looks like it would be 48 oz

OpenStudy (hunus):

Ohh, sorry. I read your problem as oz / dollar not dollar per oz. Yes \[\frac{2.96}{48} = 0.06167 < \frac{3.96}{64} = 0.061875\] 48oz has the lower unit price